OpenStudy (anonymous):

Find an integer, x, such that 5, 10, and x represent the lengths of the sides of an acute triangle.

OpenStudy (anonymous):

First x must satisfy 10-5 < x < 5+10 5 < x < 15 To have a triangle

OpenStudy (anonymous):

Now you need an acute triangle, so how do you proceed?

OpenStudy (anonymous):

If you take x near 5, then the angel opposite to 10 will be obtus

OpenStudy (anonymous):

If you take x such that x^2 = 5^2 + 10^2=125, then the angle opposite to x will be 90 degrees

OpenStudy (anonymous):

sqrt(125)=11.1803, try take x =11 and show that your triangle will be acute

OpenStudy (anonymous):

Did you understand?

OpenStudy (anonymous):

Yes, I'm beginning to understand now. Thanks for the help!